OpenStudy (jojo12g):

What value of m makes the statement true? –(m + 15) = –18 + (–15) A. –33 B. –18 C. 3 D. 18

rebeccaxhawaii (rebeccaxhawaii):

none

OpenStudy (anonymous):

first simplify both sided: -m - 15 = -18 - 15 -m - 15 = -33 then isolate the variable: -m = -18 m = 18 D

OpenStudy (jojo12g):

thats no an answer

OpenStudy (anonymous):

That is an answer